OpenStudy (anonymous):

I award metals! I have a question and i have the answers for it but I just need someone to double check them for me please.

OpenStudy (kc_kennylau):

what's your question as well as your answer? :)

OpenStudy (anonymous):

Which of the following statements have the same result? Explain each step in solving each one. I.f(3) when f(x) = 2x + 2 II.f-1(4) when f(x) = (3x-4)/5 III.y + 10 = 2y + 1

OpenStudy (anonymous):

my answer is: I. f(3)=2x+2 f(3)=2(3)+2 f(3)=6+2 f(3)=8 f(x)=8 II.F(X)=(3x-4)/5 f(x)=[3(4)-4]/5 f(x)=(12-4)/5 f(x)=8/5 f(x)=5 times 8/5 + 4 f(x)=8 + 4 f(x)=12 III.y+10=2y+1

OpenStudy (anonymous):

For the last one I got 9. But then none of them are the same so I want to know if I am right about htem being different or what.

OpenStudy (kc_kennylau):

(II) is asking you inverse?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I don't know if I did II right. Inverse always gets me

OpenStudy (kc_kennylau):

Since I got no time, I'll not guide you to the answer. I. f(3)=2(3)+2=8 II. f(x)=(3x-4)/5 x=[3f^-1(x)-4]/5 5x=3f^-1(x)-4 5x+4=3f^-1(x) f^-1(x)=(5x+4)/3 f^-1(4)=[5(4)+4]/3=8 III.y+10=2y+1 --> y=9 Therefore I and II.

OpenStudy (anonymous):

Oh okay! Thank you!! That makes more sense than what I did. Thanks again!

OpenStudy (kc_kennylau):

no problem :)